### v1

@@ Abstract
The study of $t$+$t$ cluster states in 6He provides valuable insights into exotic nuclear structures and the behavior of fermionic cluster systems. This study shows rich cluster resonant state structures above the threshold, identified by experimental reconstruction and theoretical calculations. The excitation energy spectrum above the $t$+$t$ threshold in 6He is measured via the fragmentation excitation process during the breakup reaction of 9Li on a 208Pb target at an incident energy of 32.7 MeV/nucleon. The resonant states are reconstructed from the final state coincident particles $t$+$t$ using the invariant mass method, while the non-resonant background is estimated using the event mixing method. The two new states of energy level peaks at $17.016\pm 0.002$ and $19.4\pm 0.6$ MeV are observed in addition to the previously observed energy level peaks at $13.9\pm 0.3$ and $15.0\pm 0.3$ MeV. Microscopic cluster model calculations exploring the $t+t$ resonance states in ${}^{6}\mathrm{He}$ yield theoretical energy spectra which are then compared with the current experimental results. The calculated reduced width amplitudes (RWA) of the $t+t$ channels further confirm the clustering structure of the identified $t+t$ resonance states.
@@ I Introduction
The study of nuclear cluster structures expands our understanding of nuclear physics by revealing complex interactions and collective behaviors that go beyond traditional nuclear models [1, 2, 3, 4, 5, 6]. Experimental and theoretical advances [7, 8, 9, 10, 11] explore cluster dynamics, while applications in relativistic collisions [12, 13, 14, 15] highlight their broader relevance. In the NuPECC Long Range Plan 2024 [16], it states that tritium and 3He clustering are expected to supersede the 4He clustering in neutron-rich or neutron-deficient nuclei, respectively. The exploration of these fermionic cluster systems, along with their potential to form quasi-molecular states upon adding further nucleons, is of significant interest and will pose a major challenge to nuclear physics in the coming decade.
@@
6He is a neutron-rich isotope and has an exotic nuclear structure. It is well known for its “halo” nature, where two neutrons orbit a compact core of 4He. Studying the $t$+$t$ cluster states offers insights into how such loosely-bound and neutron-rich systems form and interact. A long-standing question regarding the distinct nature of the $t$-cluster compared to the $\alpha$-cluster and its role in the formation of nuclei containing $t$-clusters [7, 17, 18, 19, 20]. While $\alpha$-clusters exhibit bosonic behavior, tritons are fermionic, making the contribution of $t$-clusters to nuclear formation a fascinating problem. Elucidating the differences between $t$-clusters and $\alpha$-clusters, offering new insights into nuclear structure. Investigating the structures of $t+\alpha$ [22, 21], $t+2\alpha$ [20, 23, 24, 25], $t+\alpha+n+n$ [26], $t+^{6}$e [7, 27, 28], $t+t$ [29], and $t+t+t$ [30], thereby advancing the cluster formation knowledge, which has garnered significant attention in both experimental and theoretical research over the past decade. Understanding the formation mechanisms of neutron-rich clusters in 6He, with the $t$-cluster playing a pivotal role. Exploring the fermionic nature of tritons and their impact on nuclear formation, addressing gaps in the study of non-$\alpha$ cluster states.
Discrepancies are evident in the reported energy level distributions of 6He across various sources; while the compilation by Ajzemberg-Selove [31] identifies only one low-lying state, Tilley et al. [32] compilation reveals two states capable of decaying into an $\alpha$ particle and two neutrons. Above the decay threshold of 12.31 MeV for 6He* states into $t+t$ clusters, the number of levels and their energies differ between the two compilations. Additional experiments, such as the observation of a 6He* state at 18.0 MeV in the 6Li(7Li, 7Be)6He [33] reaction has yielded findings not included in the aforementioned compilations. The states in the 7Li(n, d)6He [34] reaction have found the evidence of excited states at 13.6, 15.4, and 17.7 MeV above the threshold. Resonances at 15.0, 18.0, and 25.5 MeV above the $t+t$ threshold have been studied via the 6Li(7Li, 7Be t)3H reaction [35]. The study of 6He is particularly intriguing due to its behavior as an $\alpha$-particle core with a halo of two neutrons at low excitation energies, contrasting with its structure as two $t+t$ clusters at higher energies. Motivated by these discrepancies, Povoroznyk et al. [29], resolved to investigate further three-body reactions, specifically 3H($\alpha$, $tt$)1H. Their high precision observed energy levels of 6He* are 14.0$\pm$0.4, 16.1$\pm$0.4, and 18.3$\pm$0.2 MeV. with $\Gamma$ widths of 0.7$\pm$0.3, 0.8$\pm$0.4, and 1.1$\pm$0.3 MeV respectively. Nevertheless, discrepancies still exist and have not been fully clarified. The current experimental and theoretical research has made efforts to study the resonant states of $t+t$ cluster structures in 6He, and the first observation of 17.016 MeV and 19.4 MeV states has been achieved, illustrating the interesting information of energy level structures above the threshold.
@@
In the realm of theoretical research, the study of the cluster structure of the neutron-rich and unstable 6He nucleus has long captivated nuclear physicists. Initially, the intriguing ”neutron halo” in light nuclei near the drip line, particularly 6He, motivated the development of sophisticated models [36, 37]. The extended three-cluster model, which better represents the alpha particle’s internal dynamics, emphasizes core breakup effects and the role of $t+t$ clustering [38, 39, 40]. Advanced computational approaches, such as the microscopic multicluster model and resonating group method, have enhanced our understanding of the structure and dynamics of 6He, providing deeper insights into neutron-rich nuclei near the drip line [41]. While prior research on the 6He nucleus has predominantly focused on its ground and low-lying states, particularly its neutron halo structure, the theoretical investigation of resonant states above the $t+t$ threshold has been relatively underexplored. In this study, we extend the application of the Generator Coordinate Method (GCM) to investigate the $t+t$ cluster states of 6He in excitation regimes above the threshold, providing new insights into the behavior of the nucleus in these higher-energy states. GCM is a powerful tool in nuclear physics, widely used to describe complex nuclear systems by approximating collective motion through a set of quantum states [42, 43, 44]. This method constructs a wave function by integrating over a continuous set of basis states, known as generator coordinates, which correspond to different nuclear configurations. By incorporating collective degrees of freedom such as vibrations and rotations, the GCM is particularly well-suited for studying the structure and reactions of nuclei.
@@ II Experiment
Figure 1: (color online) The $E_{x}$ spectrum, reconstructed from the $t$+$t$ coincident fragments, is fitted with Breit-Wigner-shaped resonance functions (red line) together with the term of nonresonant contribution (pink dashed line marked as EM), taking the resolution (dE) and acceptance efficiency into account.
@@
As shown in FIG.  1 (a) the detectors’ spatial acceptance is $42\%$ at 13.5 MeV of excitation energy. In FIG.  1 (b), the simulation results indicated a resolution of approximately 0.4 MeV at an excited energy of 14.0 MeV, increasing to about 0.8 MeV at an excited energy of 20.0 MeV. This trend indicates that higher precision can be achieved at excitation energies near the $t+t$ threshold. The energy resolution curve is used to define bins with unequal spacing (Non-Uniform Binning), allowing for a statistical analysis of the excitation energy spectrum. Using an energy resolution curve to define bins with unequal spacing for the analysis of excitation energy spectra offers key advantages [52, 53]. This approach enhances accuracy by matching bin widths to the detector’s energy resolution, improves the signal-to-noise ratio in low-resolution regions by reducing statistical fluctuations to enhance data clarity, and optimizes data representation, ensuring efficient use of statistics without losing key spectral details.
@@
Several reaction channels are possible in the breakup reaction of ${}^{9}\text{Li}$ on a ${}^{208}\text{Pb}$ target, which leads to the production of an excited state of ${}^{6}\text{He}$. One potential reaction channel is the direct breakup, where the ${}^{9}\text{Li}$ nucleus collides with ${}^{208}\text{Pb}$ and breaks up into ${}^{6}\text{He}$ and other light fragments. Typical reaction paths include ${}^{208}\text{Pb}(^{9}\text{Li},^{6}\text{He}+t)$, ${}^{208}\text{Pb}(^{9}\text{Li},^{6}\text{He}+n+d)$, or ${}^{208}\text{Pb}(^{9}\text{Li},^{6}\text{He}+2n+p)$. Another possible channel is the coupled channel reaction, where interactions between the heavy ${}^{208}\text{Pb}$ nucleus and ${}^{9}\text{Li}$ affect the breakup process, leading to the production of an excited ${}^{6}\text{He}$. Additionally, transfer reactions may occur, where nucleons from ${}^{9}\text{Li}$ are transferred to the ${}^{208}\text{Pb}$ nucleus, leaving ${}^{6}\text{He}$ as an emission product. The reaction channel ${}^{208}\text{Pb}(^{9}\text{Li},^{6}\text{He}+t)$ suggests the decay of ${}^{9}\text{Li}$ into three tritons, with ${}^{6}\text{He}\rightarrow t+t$. However, due to the limited detection efficiency and low reaction cross-section for ${}^{9}\text{Li}$ decay into three tritons in the current experimental setup, no coincident events involving three tritons have been observed. It should be noted that the two tritons selected for reconstructing the excitation spectrum of ${}^{6}\text{He}$ might randomly be two of the three tritons from the decay of ${}^{9}\text{Li}$, contributing to the non-resonant background of ${}^{6}\text{He}$.
Understanding the excitation energy levels and the widths of excited states in light nuclei is essential for validating nuclear models and advancing research in astrophysics [29]. To analyze this, we employ a composition function of double Breit-Wigner (BW) resonance function incorporating a background term, which is convoluted with the detection resolution, while also accounting for the acceptance efficiency. This method is applied to fit the first two peaks in the $E_{x}$ spectrum, as shown in FIG.  1(a). We use double BW functions overlapped on the fitted line shapes from the first two peaks to fit the third and fourth peaks. The background contribution from non-resonant events is estimated using the ”event-mixing” (EM) technique [55, 7]. This approach, combining resonance modelling with an event-mixing technique for background estimation, ensures a more accurate description of the excitation spectrum.
Table 1: Experimental excitation energy $E_{x}$ and width $\Gamma$, fitted by Least $\chi^{2}$ method.
@@
1 13.93$\pm$0.26 0.89$\pm$0.94 0.61/2
2 14.95$\pm$0.30 0.15$\pm$1.21 0.61/2
3 17.016$\pm$0.002 0.0007$\pm$0.0039 0.04/1
4 19.44$\pm$0.56 1.57$\pm$1.64 0.04/1
The results of the Least $\chi^{2}$ fit are presented in TABLE I. The energy levels fit demonstrates a comparable precision to the study conducted by O. M. Povoroznyk et al., which provided precise measurements of the energy levels and widths of the ${}^{6}\text{He}$ state [29], surpassing earlier research [31, 32, 33, 34, 35] that often reported broader and less-defined energy levels. The first energy level shows strong agreement with the findings of M. Povoroznyk et al., while the second level at 15.0 MeV aligns with earlier studies. Notably, the third energy level, observed at 17.0 MeV with high precision, is reported here for the first time and features a very narrow width. Similarly, the 19.4 MeV level, with a width of 1.6 MeV, is also newly measured. In contrast, the levels at 16.1 and 18.3 MeV identified by Povoroznyk et al. are not observed in this work. Overall, this study achieves greater precision for the lower three energy levels, while Povoroznyk et al. demonstrate superior accuracy at higher energies. The width of the first energy level is consistent with Povoroznyk et al., though with a larger error margin. Additionally, FIG. 1(a) suggests a potential peak near 22 MeV.
@@ III Framework of GCM and RWA calculation of 6He
with $a_{1}={1.01}~{}\mathrm{fm}$, $a_{2}={1.8}~{}\mathrm{fm}$, $v_{1}={61.14}~{}\mathrm{MeV}$, $v_{2}={-60.65}~{}\mathrm{MeV}$, $W=1-M$, $M=0.6$ and $B=H=0.125$. The spin-orbit term is taken as the G3RS potential [57, 58]
@@
$\frac{\delta(r-a)}{r^{2}}\left[Y_{l}(\hat{\bm{r}})\otimes\left[\Phi_{C_{1}}^{j_{1}\pi_{1}}\otimes\Phi_{C_{2}}^{j_{2}\pi_{2}}\right]_{j_{12}}\right]_{JM}\middle|\Psi^{J\pi}_{M}\right>.$
Figure 2: (color online) The energy levels of 6He, calculated using the GCM, are provided for the 0+, 2+, 1-, 2-, and 3- states. The energy spectrum obtained from the $t+t$ cluster model is displayed in the left columns. In the middle columns, we present the results from coupling the $t+t$ configuration with the shell-model-like $\alpha+n+n$ configurations and states exhibiting significant $t+t$ structure are highlighted in red. In the middle columns, the levels depicted in black represent data from experiments, including this work and previous studies. The black dashed line indicates the potential existence of a resonance state at an excitation energy of approximately 22 MeV.
@@ IV Discussions
The energy levels above the $t+t$ threshold exhibit a rich resonance structure, including two states reported for the first time in this study. However, the levels of 16.1 and 18.3 MeV, identified by Povoroznyk et al., are not observed in this work. Both observations are performed with high precision but show inconsistency. These discrepancies highlight that, despite the rich resonance structure above the $t+t$ threshold, there is no consistent agreement across all experimental results regarding the excitation energy levels of these states. Furthermore, current measurements, including previous studies, have been unable to accurately determine the spin and parity of the excited states due to experimental accuracy and statistical data limitations. Povoroznyk et al. employed the three-body reaction pathway ${}^{3}H(\alpha,tt)^{1}H$, whereas the current experiment is based on the fragmentation (6He) excitation process during the breakup reaction of 9Li on a 208Pb target. Due to the effects of incident particle energy, angular momentum matching, nuclear barrier penetration, and selection rules, different reaction mechanisms may lead to variations in the excitation efficiency of specific energy levels. The GCM calculations (FIG.  2) suggest that the t+t resonance states of 6He may exhibit complex quantum structures, yet the absence of spin-parity measurements in the experiment limits a complete characterization of these states. The 16.1 MeV and 18.3 MeV states observed by Povoroznyk et al. may correspond to specific quantum states that were not effectively excited or identified in the current experiment’s reaction mechanism or detection setup. Future investigations should focus on spin-parity measurements, higher-statistics experiments, and multi-reaction channel comparisons to further clarify the resonance structure of 6He.
@@
To address this, we perform GCM calculations to investigate the quantum state characteristics of the excitation energy spectrum of the $t+t$ cluster structure in 6He. By employing the microscopic cluster model and effective nucleon-nucleon interactions, the theoretical results should be independent of any specific reaction channel or detection method. On the other hand, the GCM calculations suggest the spin and parity of various $t+t$ resonance states, which are crucial for unambiguously identifying these states but have not yet been provided by previous experiments. These calculations aim to provide guidance and recommendations for more precise measurements in future experiments.
Figure 3: The calculated RWA of $t+t$ configuration in the ground state and $0^{+}$ resonance states with prominent $t+t$ structure.
@@
Figure 4: The calculated RWA of $t+t$ configuration in the first excited state and $2^{+}$ resonance states with prominent $t+t$ structure.
@@
Figure 5: The calculated RWA of $t+t$ configuration in $1^{-}$ resonance states with prominent $t+t$ structure.
Figure 6: The calculated RWA of $t+t$ configuration in $2^{-}$ resonance states with prominent $t+t$ structure.
Figure 7: The calculated RWA of $t+t$ configuration in $3^{-}$ resonance states with prominent $t+t$ structure.
@@
For determining the $t+t$ resonances of ${}^{6}\mathrm{He}$ above the threshold, we perform the GCM calculation with a $t+t$ cluster model. The radius constraint method (RCM) [63] is applied to remove the continuum components from the obtained resonance states. However, we found that the energies of these resonances of $t+t$ show a strong dependence on the radius cutoff $R_{\mathrm{cut}}$. In the present calculation, we adopt $R_{\mathrm{cut}}=7.0~{}\mathrm{fm}$ with which most of the resonance states vary slowly and meanwhile, continuum components with too large radius can be efficiently eliminated. Besides, it is difficult to describe well the ground state $0_{1}^{+}$ and the first excited state $2^{+}_{1}$ of ${}^{6}\mathrm{He}$ considering merely $t+t$ configuration, as a result that these states consist of significant components of $\alpha+n+n$ clustering configurations. Consequently, we lower the energies of the ground and the first excited states by adding a few $\alpha+n+n$ configurations, in which the neutrons are close to the $\alpha$ cluster (shell-model-like) to avoid inducing continuum states.
@@
The energy spectra are shown in FIG.  2. On the left is the energy spectrum obtained by the $t+t$ cluster model. Despite the too-high ground state and first excited state, several states above the $t+t$ threshold show good agreement with the experiment results. In the middle columns, we display the result obtained by coupling the $t+t$ and shell-model-like $\alpha+n+n$ configurations. The introduction of $\alpha+n+n$ configurations significantly lowers the ground state energy as well as the first excited state energy, making them consistent with the experiment values. More importantly, we notice that some of the states obtained from $t+t$ cluster model are quite insensitive to the introduction of $\alpha+n+n$ configurations, especially for lower states. We speculate that these states should be well-developed $t+t$ resonance states, or contain considerable $t+t$ structure components. By evaluating the $t+t$ RWA for states in both the calculated spectra, we recognize the $t+t$ resonance states in the results obtained by $(t+t)\oplus(\alpha+n+n)$ model (red lines), which exhibit nearly identical RWA curves with the corresponding states obtained by $t+t$ model.

### v2

@@ Abstract
The study of tt+$t$ cluster states in 6He provides valuable insights into exotic nuclear structures and the behavior of fermionic cluster systems. This study shows rich cluster resonant state structures above the threshold, identified by experimental reconstruction and theoretical calculations. The excitation energy spectrum above the tt+$t$ threshold in 6He is measured via the fragmentation excitation process during the breakup reaction of 9Li on a 208Pb target at an incident energy of 32.7 MeV/nucleon. The resonant states are reconstructed from the final state coincident particles tt+$t$ using the invariant mass method, while the non-resonant background is estimated using the event mixing method. Two states of energy level peaks at $13.9\pm 0.3$ and $15.0\pm 0.3$ MeV are observed. Microscopic cluster model calculations exploring the $t+t$ resonant states in $^{6}\mathrm{He}$ yield theoretical energy spectra which are then compared with the current experimental results. The calculated reduced width amplitudes (RWA) of the $t+t$ channels further confirm the clustering structure of the identified $t+t$ resonant states.
@@ I Introduction
The study of nuclear cluster structures expands our understanding of nuclear physics by revealing complex interactions and collective behaviors that go beyond traditional nuclear models [1, 2, 3, 4, 5, 6]. Experimental and theoretical advances [7, 8, 9, 10, 11] explore cluster dynamics, while applications in relativistic collisions [12, 13, 14, 15, 16] highlight their broader relevance. In the NuPECC Long Range Plan 2024 [17], it states that tritium and 3He clustering are expected to supersede the 4He clustering in neutron-rich or neutron-deficient nuclei, respectively. The exploration of these fermionic cluster systems, along with their potential to form quasi-molecular states upon adding further nucleons, is of significant interest and will pose a major challenge to nuclear physics in the coming decade. Triton clustering has been investigated for several decades [18, 19, 20] and has continued to develop up to the present day [21, 22].
@@
6He is a neutron-rich isotope and has an exotic nuclear structure. It is well known for its “halo” nature, where two neutrons orbit a compact core of 4He. Studying the tt+$t$ cluster states offers insights into how such loosely-bound and neutron-rich systems form and interact. A long-standing question regarding the distinct nature of the $t$-cluster compared to the $\alpha$-cluster and its role in the formation of nuclei containing $t$-clusters [7, 23, 24, 25, 26]. While $\alpha$-clusters exhibit bosonic behavior, tritons are fermionic, making the contribution of $t$-clusters to nuclear formation a fascinating problem. Elucidating the differences between $t$-clusters and $\alpha$-clusters is expected to offer new insights into nuclear structure. Investigating the structures of $t+\alpha$ [28, 27], $t+2\alpha$ [26, 29, 30, 31], $t+\alpha+n+n$ [32], $t+^{6}$He [7, 33, 34], $t+t$ [35], and $t+t+t$ [36], thereby advancing the cluster formation knowledge, which has garnered significant attention in both experimental and theoretical research over the past decade. Understanding the formation mechanisms of neutron-rich clusters in 6He, particularly the role of the $t$ cluster, remains an important open problem. Exploring the fermionic nature of tritons and their impact on nuclear formation, addressing gaps in the study of non-$\alpha$ cluster states.
[…]
@@
In the realm of theoretical research, the study of the cluster structure of the neutron-rich and unstable 6He nucleus has long captivated nuclear physicists. Initially, the intriguing ”neutron halo” in light nuclei near the drip line, particularly 6He, motivated the development of sophisticated models [42, 43]. The extended three-cluster model, which better represents the alpha particle’s internal dynamics, emphasizes core breakup effects and the role of $t+t$ clustering [44, 45, 46]. Advanced computational approaches, such as the microscopic multicluster model and resonating group method, have enhanced our understanding of the structure and dynamics of 6He, providing deeper insights into neutron-rich nuclei near the drip line [47]. While prior research on the 6He nucleus has predominantly focused on its ground and low-lying states, particularly its neutron halo structure, the theoretical investigation of resonant states above the $t+t$ threshold has been relatively underexplored. In this study, we extend the application of the Generator Coordinate Method (GCM) to investigate the $t+t$ cluster states of 6He in excitation regimes above the threshold, providing new insights into the behavior of the nucleus in these higher-energy states. GCM is a powerful tool in nuclear physics, widely used to describe complex nuclear systems by approximating collective motion through a set of quantum states [48, 49, 50]. This method constructs a wave function by integrating over a continuous set of basis states, known as generator coordinates, which correspond to different nuclear configurations. By incorporating collective degrees of freedom such as vibrations and rotations, the GCM is particularly well-suited for studying the structure and reactions of nuclei. About forty years ago, various ${{}^{3}\mathrm{He}}+{{}^{3}\mathrm{He}}$ scattering properties, especially the spin-orbit coupling effects, were analyzed with the system wave functions obtained by GCM [51]. As a mirror system, exploring $t+t$ clustering states within the GCM framework is indispensable.
@@ II Experiment
Figure 1: (Color online) Schematic view of the detector setup [52].
Figure 2: (color online) The $E_{x}$ spectrum, reconstructed from the tt+$t$ coincident fragments, is fitted with Breit-Wigner-shaped resonance functions (red line) together with the term of nonresonant contribution (pink dashed line marked as EM), taking the resolution (dE) and acceptance efficiency into account. The curve of the acceptance efficiency is scaled by a factor of 50 for better visibility alongside the event counts.
@@
As shown in FIG.  2 (top) the detectors’ spatial acceptance is $\sim 43\%$ at 13.5 MeV of excitation energy. In FIG.  2 (bottom), the simulation results indicated a mass resolution of approximately 0.4 MeV at an excited energy of 14.0 MeV, increasing to about 0.8 MeV at an excited energy of 20.0 MeV. This trend indicates that higher precision can be achieved at excitation energies near the $t+t$ threshold. The energy resolution curve is used to define bins with unequal spacing (Non-Uniform Binning), allowing for a statistical analysis of the excitation energy spectrum. Using an energy resolution curve to define bins with unequal spacing for the analysis of excitation energy spectra offers key advantages [59, 60]. This approach enhances accuracy by matching bin widths to the detector’s energy resolution, improves the signal-to-noise ratio in low-resolution regions by reducing statistical fluctuations to enhance data clarity, and optimizes data representation, ensuring efficient use of statistics without losing key spectral details.
@@
Several reaction channels are possible in the breakup reaction of $^{9}\text{Li}$ on a $^{208}\text{Pb}$ target, which leads to the production of an excited state of $^{6}\text{He}$. One potential reaction channel is the direct breakup, where the $^{9}\text{Li}$ nucleus collides with $^{208}\text{Pb}$ and breaks up into $^{6}\text{He}$ and other light fragments. Typical reaction paths include $^{208}\text{Pb}(^{9}\text{Li},^{6}\text{He}+t)$, $^{208}\text{Pb}(^{9}\text{Li},^{6}\text{He}+n+d)$, or $^{208}\text{Pb}(^{9}\text{Li},^{6}\text{He}+2n+p)$. Another possible channel is the coupled channel reaction, where interactions between the heavy $^{208}\text{Pb}$ nucleus and $^{9}\text{Li}$ affect the breakup process, leading to the production of an excited $^{6}\text{He}$. Additionally, transfer reactions may occur, where nucleons from $^{9}\text{Li}$ are transferred to the $^{208}\text{Pb}$ nucleus, leaving $^{6}\text{He}$ as an emission product. The reaction channel $^{208}\text{Pb}(^{9}\text{Li},^{6}\text{He}+t)$ suggests the decay of $^{9}\text{Li}$ into three tritons, with $^{6}\text{He}\rightarrow t+t$. However, due to the limited detection efficiency and low reaction cross-section for $^{9}\text{Li}$ decay into three tritons in the current experimental setup, no coincident events involving three tritons have been observed. It should be noted that if $^{9}\text{Li}$ decays via a cascade process ($^{9}\text{Li}\rightarrow{{}^{6}\text{He}}+t$ followed by ${{}^{6}\text{He}}\rightarrow t+t$), only the two tritons originating from the $^{6}\text{He}$} decay are considered as signal. In contrast, any pair consisting of one triton from the $^{6}\text{He}$} decay and one from the initial step, as well as any pair from a direct $^{9}\text{Li}\rightarrow t+t+t$ decay, is treated as part of the non-resonant background and is estimated using the event-mixing method [62].
Understanding the excitation energy levels and the widths of excited states in light nuclei is essential for validating nuclear models and advancing research in astrophysics [35]. To analyze this, we employ a composition function of double Breit-Wigner (BW) resonance function incorporating a background term, which is convoluted with the detection resolution, while also accounting for the acceptance efficiency. This method is applied to fit the first two peaks in the $E_{x}$ spectrum, as shown in FIG.  2. We use double BW functions overlapped on the fitted line shapes from the first two peaks to profile the third and fourth peaks. However, due to an insufficient number of data points, no reliable fitting parameters could be obtained for these higher-order peaks. The background contribution from non-resonant events is estimated using the ”event-mixing” (EM) technique [62, 7]. This approach, combining resonance modeling with an event-mixing technique for background estimation, ensures a more accurate description of the excitation spectrum.
Table 1: Experimental excitation energy $E_{x}$ and width $\Gamma$, fitted by Least $\chi^{2}$ method with non-uniform binning.
@@
[…]
The results of the Least $\chi^{2}$ fit with non-uniform binning are presented in TABLE I. The uncertainties for the peak parameters are the errors obtained from the least-squares fit applied to the reconstructed invariant mass spectrum. These errors, derived from the covariance matrix during fitting, represent statistical uncertainties under the assumption that the model is correct and that data errors are known. In our fitting process, the measured peak shape is deconvolved with the estimated energy resolution to extract the intrinsic peak parameters. Moreover, fitting four peaks requires 12 parameters, plus one background parameter, totaling 13 parameters. This is indeed challenging with limited data, and appropriate binning is essential to ensure a reliable fit. We adopted non-uniform binning based on the energy resolution curve, which optimizes the signal-to-noise ratio by matching bin widths to the resolution and minimizing statistical fluctuations. This approach enhances data clarity and determines a stable profile. However, in high-excitation energy regions with wide bins, the effective number of bins is reduced, thus lowering the fit’s degrees of freedom and making fitting parameters less reliable.
The energy levels fit demonstrates a comparable precision to the study conducted by O. M. Povoroznyk et al., which provided precise measurements of the energy levels and widths of the $^{6}\text{He}$ state [35], surpassing earlier research [37, 38, 39, 40, 41] that often reported broader and less-defined energy levels. The first energy level at $13.9\pm 0.3$ MeV shows strong agreement with the findings of M. Povoroznyk et al., while the second level at $15.0\pm 0.3$ MeV aligns with earlier studies. Due to its limited statistics, two candidate peaks around 16.9 MeV and 19.5 MeV need to be further determined. In contrast, the levels at 16.1 and 18.3 MeV identified by Povoroznyk et al. are not observed in this work. This study achieves the same precision for the lower two energy levels, while Povoroznyk et al. demonstrate superior accuracy at higher energies. The width of the first energy level is consistent with Povoroznyk et al., though with a larger error margin. Overall, the fitting errors for width are large, and future precise measurements are essential.
@@ III Framework of GCM and RWA calculation of 6He
with $a_{1}={1.01}~\mathrm{fm}$, $a_{2}={1.8}~\mathrm{fm}$, $v_{1}={61.14}~\mathrm{MeV}$, $v_{2}={-60.65}~\mathrm{MeV}$, W=1MW=1-M, $M=0.6$ and $B=H=0.125$. The spin-orbit term is taken as the G3RS potential [64, 65]
@@
$\left<\frac{\delta(r-a)}{r^{2}}\left[Y_{l}(\hat{\bm{r}})\otimes\left[\Phi_{C_{1}}^{j_{1}\pi_{1}}\otimes\Phi_{C_{2}}^{j_{2}\pi_{2}}\right]_{j_{12}}\right]_{JM}\middle|\Psi^{J\pi}_{M}\right$>.
Figure 3: (color online) The energy levels of 6He, calculated using the GCM, are provided for the 0+, 2+, 1-, 2-, and 3- states. The energy spectrum obtained from the $t+t$ cluster model is displayed in the left columns. In the middle columns, we present the results from coupling the $t+t$ configuration with the shell-model-like $\alpha+n+n$ configurations and states exhibiting significant $t+t$ structure are highlighted in red. In the right columns, the levels depicted in black represent data from experiments, including this work and previous studies. The dashed lines are the candidates for further confirmation.
@@ IV Discussions
The energy levels above the $t+t$ threshold exhibit a rich resonance structure, including two states $13.9$ and $15.0$ MeV reported in this study. However, the levels of 16.1 and 18.3 MeV, identified by Povoroznyk et al., are not observed in this work. These discrepancies highlight that, despite the rich resonance structure above the $t+t$ threshold, there is no consistent agreement across all experimental results regarding the excitation energy levels of these states. Furthermore, current measurements, including previous studies, have been unable to accurately determine the spin and parity of the excited states due to experimental accuracy and statistical data limitations. Povoroznyk et al. employed the three-body reaction pathway $^{3}H(\alpha,tt)^{1}H$, whereas the current experiment is based on the fragmentation (6He) excitation process during the breakup reaction of 9Li on a 208Pb target. Due to the effects of incident particle energy, angular momentum matching, nuclear barrier penetration, and selection rules, different reaction mechanisms may lead to variations in the excitation efficiency of specific energy levels. The GCM calculations (FIG.  3) suggest that the t+t resonant states of 6He may exhibit complex quantum structures, yet the absence of spin-parity measurements in the experiment limits a complete characterization of these states. The 16.1 MeV and 18.3 MeV states observed by Povoroznyk et al. may correspond to specific quantum states that were not effectively excited or identified in the current experiment’s reaction mechanism or detection setup. Future investigations should focus on spin-parity measurements, higher-statistics experiments, and multi-reaction channel comparisons to further clarify the resonance structure of 6He.
@@
To address this, we perform GCM calculations to investigate the quantum state characteristics of the excitation energy spectrum of the $t+t$ cluster structure in 6He. By employing the microscopic cluster model and effective nucleon-nucleon interactions, the theoretical results should be independent of any specific reaction channel or detection method. On the other hand, the GCM calculations suggest the spin and parity of various $t+t$ resonant states, which are crucial for unambiguously identifying these states but have not yet been provided by previous experiments. These calculations aim to provide guidance and recommendations for more precise measurements in future experiments.
@@
Figure 4: The calculated RWA of $t+t$ configuration in the ground state, the first excited state, and positive-parity resonant states with prominent $t+t$ structure.
@@
Figure 5: The calculated RWA of $t+t$ configuration in negative-parity resonant states with prominent $t+t$ structure.
@@
For determining the $t+t$ resonances of $^{6}\mathrm{He}$ above the threshold, we perform the GCM calculation with a $t+t$ cluster model. The radius constraint method (RCM) [70] is applied to remove the continuum components from the obtained resonant states. However, we found that the energies of these resonances of $t+t$ show a strong dependence on the radius cutoff $R_{\mathrm{cut}}$. In the present calculation, we adopt $R_{\mathrm{cut}}=5.0~\mathrm{fm}$. Besides, it is difficult to describe well the ground state $0_{1}^{+}$ and the first excited state $2^{+}_{1}$ of $^{6}\mathrm{He}$ considering merely $t+t$ configuration, as a result that these states consist of significant components of $\alpha+n+n$ clustering configurations. Consequently, we lower the energies of the ground and the first excited states by adding a few $\alpha+n+n$ configurations, in which the neutrons are close to the $\alpha$ cluster (shell-model-like) to avoid inducing continuum states.
@@
The energy spectra are shown in FIG. 3. On the left is the energy spectrum obtained by the $t+t$ cluster model. Despite the too-high ground state and first excited state, several states above the $t+t$ threshold show good agreement with the experiment results. In the middle columns, we display the result obtained by coupling the $t+t$ and shell-model-like $\alpha+n+n$ configurations. The introduction of $\alpha+n+n$ configurations significantly lowers the ground state energy as well as the first excited state energy, making them consistent with the experimental values. More importantly, we notice that some of the states obtained from the $t+t$ cluster model are quite insensitive to the introduction of $\alpha+n+n$ configurations. We speculate that these states should be well-developed $t+t$ resonance states, or contain considerable $t+t$ structure components. By evaluating the $t+t$ RWA for states in both the calculated spectra, we recognize the $t+t$ resonance states in the results obtained by the $(t+t)\oplus(\alpha+n+n)$ model (red lines), which exhibit nearly identical RWA curves with the corresponding states obtained by $t+t$ model.